how many 2 inch square blocks can fit inside a rectangle with the measurements of 10inches on both sides and 20 inches in on the top and bottom. the answer is 50 but how do u get the answer
Answer by mananth(16949)

The volume of box = L*W*H
Volume = 20*20*10= 400 in^3
..
Volume of small cube = 2*2*2= 8 in^3
..
Number of cubes = volume of bigger cuboid/volume of smaller cuboid.
Number = 400/8= 50
...
m.ananth@hotmail.ca